# Spectraformer: A Unified Random Feature Framework for Transformer

### v2

@@ A.3 Technical Details of Weight Approximators
SSR (Spherical-Radial Rules) [55] transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
$\boldsymbol{W}_{\text{SSR}}=\boldsymbol{\vartheta}\otimes\begin{bmatrix}(\boldsymbol{QV})^{\top}\\ -(\boldsymbol{Q}\boldsymbol{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
where $\boldsymbol{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\boldsymbol{V}=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\boldsymbol{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\boldsymbol{Q}$ (a random orthogonal matrix).
@@ A.4 Technical Details of Weight Learners
We give a non-mathematical definition of weight matrix learner in Section 3.1: weight matrix learner is any function which parameterize distributions. This allows for the discovery and future experimentation with other potential weight matrix learners in Spectraformer and random feature method in general. Weight matrix learner comes from a long line of work in kernel methods, namely, kernel learning via random features. Typically, a component function $f$ is chosen and a weight matrix $\boldsymbol{W}$ is produced via kernel approximation, providing us with the feature map $\phi$. $\phi$ can then be combined with input $\boldsymbol{x}$ and weight $\boldsymbol{v}$ to fit the learning objective. Kernel learning, however, imposes an additional objective: learning the weight matrix $\boldsymbol{W}$.
@@
The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods [7]. Two-stage methods (e.g., [56], [23], [57]) solves the dual objective separately: $\boldsymbol{W}$ is typically first learned via a kernel alignment scheme [58], then we solve for $\boldsymbol{v}$. One-stage methods (e.g., [14], [22]), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\boldsymbol{W}$ is parameterized then $\boldsymbol{v}$ is solved typically. It is not plausible to apply kernel alignment [58] on Transformer architectures. Therefore, only one-stage methods can be considered in the context of weight matrix learner. Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting [29], then adapted into the Transformer setting [14]. DGM (Deep Generative Model) has also been effective in modeling distributions [59]. It too has been adapted by [14] for the Transformer setting. Due to the superior performance of FastFoodL in [14], we have only experimented with this technique. We now explore them.
@@
FastFoodL (see Equation 17) is proposed to be learnable by [29] with two variations: FSARD (scaling matrix $\boldsymbol{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\boldsymbol{G}$ and $\boldsymbol{B}$). It is further adapted by [14] to make $\boldsymbol{S}$ or $\boldsymbol{S},\boldsymbol{G},\boldsymbol{B}$ learnable parameters. In our experiments, we maintain the set up by [14].
@@
GMM (Gaussian Mixture Model) [23, 24] is a known universal approximator of probability distribution. Specifically, we can consider $\boldsymbol{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$. $p(\boldsymbol{\omega})$ can then be approximated as $p(\boldsymbol{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ [14]. Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as [14]:
$\boldsymbol{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ (29)
@@
Another formulation of GMM is proposed in [29] in combination with FastFood, albeit it extends beyond constructing the weight matrix.
@@
DGM (Deep Generative Model) [59] allows for a generic distribution modeling using neural network. DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\boldsymbol{\omega}_{m}=g(\boldsymbol{n}_{m})$ via transforming from a noise distribution $\boldsymbol{n}_{m}\sim p_{0}(.)$.
@@
In addition to the techniques discussed above, there are also some other prominent one-stage methods (e.g., INN, BBQ). INN (Invertible Neural Network) [60] is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution. BBQ (Black Box Quantile) [22] follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
@@ A.5 Technical Details of Component Functions
Figure 5: Spectraformer component functions. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work.
@@
CDERF (Complex dense exponential random feature) is given in Equation 30, where $\boldsymbol{\omega},\boldsymbol{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices. The parameters and constraints of different CDERF functions are specified in Table 4:
$f_{DE}^{(k)}(\boldsymbol{\omega},\boldsymbol{x})=Real(D\exp(\boldsymbol{\omega}^{T}\boldsymbol{A}\boldsymbol{\omega}+\boldsymbol{\omega}^{T}\boldsymbol{B}^{(k)}\boldsymbol{x}+\boldsymbol{x}^{T}\boldsymbol{C}^{(k)}\boldsymbol{x}))$ (30)
@@
$\boldsymbol{A}$ $\boldsymbol{B}^{(1)}$ $\boldsymbol{B}^{(2)}$ $\boldsymbol{C}^{(1)}$ $\boldsymbol{C}^{(2)}$ $D$ $\boldsymbol{x}$ $\boldsymbol{y}$
@@
SADERF [13] $A_{GE}\boldsymbol{I}_{d}$ $B_{GE}\boldsymbol{I}_{d}$ $C_{GE}\boldsymbol{I}_{d}$ $D_{GE}$ $\Psi\boldsymbol{x}$ $\Psi^{-1}\boldsymbol{y}$
GERF [12] $A_{GE}\boldsymbol{I}_{d}$ $B_{GE}^{(1)}\boldsymbol{I}_{d}$ $B_{GE}^{(2)}\boldsymbol{I}_{d}$ $C_{GE}\boldsymbol{I}_{d}$ $D_{GE}$ $\boldsymbol{x}$ $\boldsymbol{y}$
TrigRF [19] $0\boldsymbol{I}_{d}$ $i\boldsymbol{I}_{d}$ $-i\boldsymbol{I}_{d}$ $0\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
PosRF [8] $0\boldsymbol{I}_{d}$ $1\boldsymbol{I}_{d}$ $-1\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
OPRF [12] $0\boldsymbol{I}_{d}$ $1\boldsymbol{I}_{d}$ $-1\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
TrigRF (Trigonometric RF) [19] is the base RFF implementation.
$\begin{array}[]{c}f_{TrigRF}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=\sqrt{2}\cos(\boldsymbol{\omega}^{T}\boldsymbol{x}+\boldsymbol{b})\\ f_{TrigRF}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=\sqrt{2}\sin(\boldsymbol{\omega}^{T}\boldsymbol{y}+\boldsymbol{b})\\ \boldsymbol{b}\sim Uniform(0,2\pi)\end{array}$ (31)
The use of the trigonometric sine and cosine functions leads to unstable behavior when the inputs have negative dimension-values. This can be further exacerbated when the values TrigRF try to approximate are close to 0 (since most values are of low significance). This causes the variance to approach infinity [8, 12]. Therefore, we do not want to use TrigRF to perform kernel approximation in attention.
PosRF (Positive RF) [8] fixes the problem of TrigRF by enforcing positive component function output in the softmax. The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0. PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 32) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 33).
$\begin{array}[]{c}f_{PosRF_{B}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(\boldsymbol{\omega}^{T}\boldsymbol{x}-\frac{||\boldsymbol{x}||^{2}}{2})\end{array}$ (32)
$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(\boldsymbol{\omega}^{T}\boldsymbol{x}-||\boldsymbol{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(-\boldsymbol{\omega}^{T}\boldsymbol{x}-||\boldsymbol{x}||^{2})\end{array}$ (33)
@@
GERF (Generalized exponential RF) [12] generalizes both TrigRF and PosRF with Equation 34.
$\begin{array}[]{c}f_{GERF}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=D\exp(A||\boldsymbol{\omega}||^{2}+B\boldsymbol{\omega}^{T}\boldsymbol{x}+\boldsymbol{C}||\boldsymbol{x}||^{2})\\ f_{GERF}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=D\exp(\boldsymbol{A}||\boldsymbol{\omega}||^{2}+sB\boldsymbol{\omega}^{T}\boldsymbol{y}+\boldsymbol{C}||\boldsymbol{y}||^{2})\\ \\ Re(1-4A)>0,\boldsymbol{B}=\sqrt{s(1-4\boldsymbol{A})}\\ \boldsymbol{C}=-(s+1)/2,D=(\sqrt[4]{1-4\boldsymbol{A}})^{d}\\ \boldsymbol{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
@@
OPRF (Optimized positive RF) [12] is the solution to the minimization of the variance of GERF. Specifically it is defined as Equation 34 with $s=+1$, $||\boldsymbol{x}+\boldsymbol{y}||^{2}>0$, and $\boldsymbol{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
$\begin{array}[]{cc}\boldsymbol{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\boldsymbol{x}+\boldsymbol{y}||^{2}+d)^{2}+8d||\boldsymbol{x}+\boldsymbol{y}||^{2}}-2||\boldsymbol{x}+\boldsymbol{y}||^{2}-d)\\ /(4||\boldsymbol{x}+\boldsymbol{y}||^{2})\end{array}$ (35)
Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$$\times$ variance reduction in estimating the Softmax compared to TrigRF [12].
DERF (Dense-exponential random features) [13] extends GERF and replace $\boldsymbol{A},\boldsymbol{B},\boldsymbol{C}$ with dense matrices. DERF is CDERF when $\boldsymbol{B},C$ are in the real instead of the complex plane. Minimizing the variance of DERF leads to two approaches: ADERF and SDERF. However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries. Therefore SADERF is proposed.
SADERF (simplified ADERF) [13] is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
$\begin{array}[]{c}f_{SADE}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=f_{GE}^{(1)}(\boldsymbol{\omega},\Psi\boldsymbol{x}),f_{SADE}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=f_{GE}^{(2)}(\boldsymbol{\omega},\Psi^{-1}\boldsymbol{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\boldsymbol{y}_{l}^{(i)})^{2}/\sum_{i}(\boldsymbol{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
DIRF (Discretely-induced random features) [12] are based on the assumption of $p(\boldsymbol{\omega})$ being a discrete distribution with $\boldsymbol{\omega}_{1},...,\boldsymbol{\omega}_{d}$ being i.i.d., $\mathbb{P}(\boldsymbol{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$. By making use of the Taylor series, we can define DIRF as:
$f_{DI}(\boldsymbol{\omega},\boldsymbol{x})=\exp(-\frac{||\boldsymbol{x}||^{2}}{2})\prod_{l=1}^{d}\boldsymbol{x}_{i}^{\boldsymbol{\omega}_{l}}(\boldsymbol{\omega}_{l}!)^{\frac{-1}{2}}p_{\boldsymbol{\omega}_{l}}^{\frac{-1}{2}}$ (37)
@@
HRF (Hybrid RF) [10] combines multiple base estimators, similar to multiple kernel learning methods [61], in order to provide the most accurate approximation in regions of interest. Choromanski et al. on solving softmax approximation using HRF. HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\boldsymbol{x},\boldsymbol{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\boldsymbol{x},\boldsymbol{y})\hat{SM}^{k}(\boldsymbol{x},\boldsymbol{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\boldsymbol{x},\boldsymbol{y}))\hat{SM}^{p+1}(\boldsymbol{x},\boldsymbol{y})$ (38)
Since HRF requires the base estimators $SM(\boldsymbol{x},\boldsymbol{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\boldsymbol{x},\boldsymbol{y})=\psi^{m}_{A}(\boldsymbol{x})^{T}\psi^{m}_{(\boldsymbol{A}^{T})^{-1}}(\boldsymbol{y})\\ \psi^{m}_{\boldsymbol{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\boldsymbol{Mu})^{2}}{2})\\ (\exp(\boldsymbol{\omega}^{T}_{i}\boldsymbol{Mu}),...,\exp(\boldsymbol{\omega}_{m}^{T}\boldsymbol{Mu}))^{T},\boldsymbol{\omega}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d})\end{array}$ (39)

### v3

@@ A.3 Technical Details of Weight Approximators
SSR (Spherical-Radial Rules) (Munkhoeva et al., 2018) transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
$\bm{W}_{\text{SSR}}=\bm{\vartheta}\otimes\begin{bmatrix}(\bm{QV})^{\top}\\ -(\bm{Q}\bm{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
where $\bm{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\bm{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\bm{Q}$ (a random orthogonal matrix).
@@ A.4 Technical Details of Weight Learners
We give a non-mathematical definition of weight matrix learner in Section 3.1: weight matrix learner is any function which parameterize distributions. This allows for the discovery and future experimentation with other potential weight matrix learners in Spectraformer and random feature method in general. Weight matrix learner comes from a long line of work in kernel methods, namely, kernel learning via random features. Typically, a component function $f$ is chosen and a weight matrix $\bm{W}$ is produced via kernel approximation, providing us with the feature map $\phi$. $\phi$ can then be combined with input $\bm{x}$ and weight $\bm{v}$ to fit the learning objective. Kernel learning, however, imposes an additional objective: learning the weight matrix $\bm{W}$.
@@
The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods (Liu et al., 2022). Two-stage methods (e.g., Yu et al. (2015); Wilson & Adams (2013); Bullins et al. (2018)) solves the dual objective separately: $\bm{W}$ is typically first learned via a kernel alignment scheme (Wang et al., 2015), then we solve for $\bm{v}$. One-stage methods (e.g., Chowdhury et al. (2022); Tompkins et al. (2019)), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\bm{W}$ is parameterized then $\bm{v}$ is solved typically. It is not plausible to apply kernel alignment (Wang et al., 2015) on Transformer architectures. Therefore, only one-stage methods can be considered in the context of weight matrix learner. Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting (Yang et al., 2015), then adapted into the Transformer setting (Chowdhury et al., 2022). DGM (Deep Generative Model) has also been effective in modeling distributions (Kingma & Welling, 2013). It too has been adapted by Chowdhury et al. (2022) for the Transformer setting. Due to the superior performance of FastFoodL with this technique, we have only experimented with this technique. We now explore them.
@@
FastFoodL (see Equation 17) is proposed to be learnable by Yang et al. (2015) with two variations: FSARD (scaling matrix $\bm{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\bm{G}$ and $\bm{B}$). It is further adapted by Chowdhury et al. (2022) to make $\bm{S}$ or $\bm{S},\bm{G},\bm{B}$ learnable parameters. In our experiments, we maintain this set up.
@@
GMM (Gaussian Mixture Model) (Wilson & Adams, 2013; Oliva et al., 2016) is a known universal approximator of probability distribution. Specifically, we can consider $\bm{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$. $p(\bm{\omega})$ can then be approximated as $p(\bm{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ (Chowdhury et al., 2022). Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as:
$\bm{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\bm{0},\bm{I})$ (29)
@@
Another formulation of GMM is proposed in Yang et al. (2015) in combination with FastFood, albeit it extends beyond constructing the weight matrix.
@@
DGM (Deep Generative Model) (Kingma & Welling, 2013) allows for a generic distribution modeling using neural network. DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\bm{\omega}_{m}=g(\bm{n}_{m})$ via transforming from a noise distribution $\bm{n}_{m}\sim p_{0}(.)$.
@@
In addition to the techniques discussed above, there are also some other prominent one-stage methods (e.g., INN, BBQ). INN (Invertible Neural Network) (Xu et al., 2021) is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution. BBQ (Black Box Quantile) (Tompkins et al., 2019) follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
@@ A.5 Technical Details of Component Functions
Figure 3: Spectraformer component functions. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work. Please refer to Appendix A.5 for references.
@@
CDERF (Complex dense exponential random feature) is given in Equation 30, where $\bm{\omega},\bm{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices. The parameters and constraints of different CDERF functions are specified in Table 7:
$f_{DE}^{(k)}(\bm{\omega},\bm{x})=Real(D\exp(\bm{\omega}^{T}\bm{A}\bm{\omega}+\bm{\omega}^{T}\bm{B}^{(k)}\bm{x}+\bm{x}^{T}\bm{C}^{(k)}\bm{x}))$ (30)
@@
$\bm{A}$ $\bm{B}^{(1)}$ $\bm{B}^{(2)}$ $\bm{C}^{(1)}$ $\bm{C}^{(2)}$ $D$ $\bm{x}$ $\bm{y}$
@@
SADERF (Likhosherstov et al., 2023) $A_{GE}\bm{I}_{d}$ $B_{GE}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\Psi\bm{x}$ $\Psi^{-1}\bm{y}$
GERF (Likhosherstov et al., 2022) $A_{GE}\bm{I}_{d}$ $B_{GE}^{(1)}\bm{I}_{d}$ $B_{GE}^{(2)}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\bm{x}$ $\bm{y}$
TrigRF (Rahimi & Recht, 2007) $0\bm{I}_{d}$ $i\bm{I}_{d}$ $-i\bm{I}_{d}$ $0\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
PosRF (Choromanski et al., 2021) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
OPRF (Likhosherstov et al., 2022) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
TrigRF (Trigonometric RF) (Rahimi & Recht, 2007) is the base RFF implementation.
$\begin{array}[]{c}f_{TrigRF}^{(1)}(\bm{\omega},\bm{x})=\sqrt{2}\cos(\bm{\omega}^{T}\bm{x}+\bm{b})\\ f_{TrigRF}^{(2)}(\bm{\omega},\bm{y})=\sqrt{2}\sin(\bm{\omega}^{T}\bm{y}+\bm{b})\\ \bm{b}\sim Uniform(0,2\pi)\end{array}$ (31)
The use of the trigonometric sine and cosine functions leads to unstable behavior when the inputs have negative dimension-values. This can be further exacerbated when the values TrigRF try to approximate are close to 0 (since most values are of low significance). This causes the variance to approach infinity (Choromanski et al., 2021; Likhosherstov et al., 2022). Therefore, we do not want to use TrigRF to perform kernel approximation in attention.
PosRF (Positive RF) (Choromanski et al., 2021) fixes the problem of TrigRF by enforcing positive component function output in the softmax. The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0. PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 32) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 33).
$\begin{array}[]{c}f_{PosRF_{B}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-\frac{||\bm{x}||^{2}}{2})\end{array}$ (32)
$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(-\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\end{array}$ (33)
@@
GERF (Generalized exponential RF) (Likhosherstov et al., 2022) generalizes both TrigRF and PosRF with Equation 34.
$\begin{array}[]{c}f_{GERF}^{(1)}(\bm{\omega},\bm{x})=D\exp(A||\bm{\omega}||^{2}+B\bm{\omega}^{T}\bm{x}+\bm{C}||\bm{x}||^{2})\\ f_{GERF}^{(2)}(\bm{\omega},\bm{y})=D\exp(\bm{A}||\bm{\omega}||^{2}+sB\bm{\omega}^{T}\bm{y}+\bm{C}||\bm{y}||^{2})\\ \\ Re(1-4A)>0,\bm{B}=\sqrt{s(1-4\bm{A})}\\ \bm{C}=-(s+1)/2,D=(\sqrt[4]{1-4\bm{A}})^{d}\\ \bm{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
@@
OPRF (Optimized positive RF) (Likhosherstov et al., 2022) is the solution to the minimization of the variance of GERF. Specifically it is defined as Equation 34 with $s=+1$, $||\bm{x}+\bm{y}||^{2}>0$, and $\bm{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
$\begin{array}[]{cc}\bm{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\bm{x}+\bm{y}||^{2}+d)^{2}+8d||\bm{x}+\bm{y}||^{2}}-2||\bm{x}+\bm{y}||^{2}-d)\\ /(4||\bm{x}+\bm{y}||^{2})\end{array}$ (35)
Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$$\times$ variance reduction in estimating the Softmax compared to TrigRF.
DERF (Dense-exponential random features) (Likhosherstov et al., 2023) extends GERF and replace $\bm{A},\bm{B},\bm{C}$ with dense matrices. DERF is CDERF when $\bm{B},C$ are in the real instead of the complex plane. Minimizing the variance of DERF leads to two approaches: ADERF and SDERF. However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries. Therefore SADERF is proposed.
SADERF (simplified ADERF) (Likhosherstov et al., 2023) is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
$\begin{array}[]{c}f_{SADE}^{(1)}(\bm{\omega},\bm{x})=f_{GE}^{(1)}(\bm{\omega},\Psi\bm{x}),f_{SADE}^{(2)}(\bm{\omega},\bm{y})=f_{GE}^{(2)}(\bm{\omega},\Psi^{-1}\bm{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\bm{y}_{l}^{(i)})^{2}/\sum_{i}(\bm{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
DIRF (Discretely-induced random features) (Likhosherstov et al., 2022) are based on the assumption of $p(\bm{\omega})$ being a discrete distribution with $\bm{\omega}_{1},...,\bm{\omega}_{d}$ being i.i.d., $\mathbb{P}(\bm{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$. By making use of the Taylor series, we can define DIRF as:
$f_{DI}(\bm{\omega},\bm{x})=\exp(-\frac{||\bm{x}||^{2}}{2})\prod_{l=1}^{d}\bm{x}_{i}^{\bm{\omega}_{l}}(\bm{\omega}_{l}!)^{\frac{-1}{2}}p_{\bm{\omega}_{l}}^{\frac{-1}{2}}$ (37)
@@
HRF (Hybrid RF) (Choromanski et al., 2022) combines multiple base estimators, similar to multiple kernel learning methods (Bach et al., 2004), in order to provide the most accurate approximation in regions of interest. Choromanski et al. on solving softmax approximation using HRF. HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\bm{x},\bm{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y})\hat{SM}^{k}(\bm{x},\bm{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y}))\hat{SM}^{p+1}(\bm{x},\bm{y})$ (38)
Since HRF requires the base estimators $SM(\bm{x},\bm{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\bm{x},\bm{y})=\psi^{m}_{A}(\bm{x})^{T}\psi^{m}_{(\bm{A}^{T})^{-1}}(\bm{y})\\ \psi^{m}_{\bm{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\bm{Mu})^{2}}{2})\\ (\exp(\bm{\omega}^{T}_{i}\bm{Mu}),...,\exp(\bm{\omega}_{m}^{T}\bm{Mu}))^{T},\bm{\omega}_{i}\sim\mathcal{N}(\bm{0},\bm{I}_{d})\end{array}$ (39)